Find the equation of the line that has a slope of -3/4 and passes through the point (1,2)
Answer by longjonsilver(2297)

equation of straight line is y=mx+c, m is the gradient, c is the y-intercept.
so, point(1,2) means x=1, y=2, so y=mx+c becomes...
2 = -3/4(1) + c
2 = -3/4 + c
c=2.75
so equation is y=-0.75x + 2.75

jon